OpenStudy (anonymous):

solve by substitution x+5y=5 4x+6y=6

OpenStudy (anonymous):

solve it on ur own

OpenStudy (anonymous):

(-1,0)

OpenStudy (anonymous):

x+5y=5 x=5-5y----------eqn1 4x+6y=6 4(5-5y)+6y=6 20-25y+6y=6 -19y=-19 y=1 then x =0

OpenStudy (anonymous):

1, 0